CRV is positive is 3/8. . A CRV has uniform distribution from -5 to 3. The probability that the i. | If X has PMF beside and Y=X, then E(Y)=1.8 : / p(r) 1/15 2/15 3/15 4/16 5/15 ) The number of different 8-bit words containing three ones and five zeros are 40G j. If PA) 0.4, P(B) 0.3 and P(AUB)C)0.42 then A, B are dependent. Two mutually-exclusive events are dependent. . The slope of the PDF for CRV couldnt be negative.

True or False With explanation please.
Answer #1

g:

The probability that the CRV is positive is

3 3-(-5) L8J08

Hence, it is true.

h:

Following table shows the calculations:

X p(x) Y=X^2 p(y)=p(x) y*p(y)
-2 1/15 4 1/15 4/15
-1 2/15 1 2/15 2/15
0 3/15 0 3/15 0
1 4/15 1 4/15 4/15
2 5/15 4 5/15 20/15
Total 2

So,

E(Y)=sum yp(y)=2

Hence, it is false.

i:

The possible number of permutations of 3 ones and five zeros so

8! 10-56

Hence, it is false.

j

If A and B are independent then

P(A or B) = P(A) + P(B)-P(A and B) = 0.4+0.3 - 0.4*0.3 = 0.58

So

P ((A or B)*) = 1-0.58 = 0.42

That is A and B are independent. Hence, it is false.

k:

False

l:

Slope of CDF cannot be negative.

False